OpenStudy (anonymous):

please help! Medal and fan Which of the following cannot be solved using the quadratic formula? A. -3x^2-x+3 =-3x^2 B. 0=x^2 C. 0=2x^2-1 D. x^2 +3x-2=4x^2-5x-1

OpenStudy (jdoe0001):

check their discriminant if the discriminant is a negative value, then it has no solution or two complex roots http://www.regentsprep.org/regents/math/algtrig/ate3/discriminant.htm

OpenStudy (anonymous):

but you need 3 values for a discriminant, a,b,and c. and B. doesnt have all 3. im confused

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

hmmm actually it should be \(\bf A. \quad -3x^2-x+3 =-3x^2\implies \cancel{ -3x^2+3x^2}-x+3 =0 \\ \quad \\ \quad y={\color{blue}{ 0}}x^2{\color{red}{ -1}}x{\color{green}{ +3}} \qquad \qquad {{\color{red}{ b}}^2 -4{\color{blue}{ a}}{\color{green}{ c}}}\)

OpenStudy (anonymous):

are you saying its a?

OpenStudy (jdoe0001):

nope, I'm saying that all have 3 components but sometimes one of them may well be 0 lemme rewrite all 4

OpenStudy (jdoe0001):

\(\bf A. \quad -3x^2-x+3 =-3x^2\implies \cancel{ -3x^2+3x^2}-x+3 =0 \\ \quad \\ \quad {\color{blue}{ 0}}x^2{\color{red}{ -1}}x{\color{green}{ +3}}=0\qquad \qquad \\ \quad \\ % --------------------- B.\quad 0=x^2\implies 0={\color{blue}{ 1}}x^2{\color{red}{ +0}}x{\color{green}{ +0}} \\ \quad \\ %---------------------- C.\quad 0=2x^2-1\implies 0={\color{blue}{ 2}}x^2{\color{red}{ +0}}x{\color{green}{ -1}} \\ \quad \\ %--------------------- D. \quad x^2 +3x-2=4x^2-5x-1 \\ \quad \\ 0=4x^2-x^2-5x-3x-1+2\implies 0={\color{blue}{ 3}}x^2{\color{red}{ -8}}x{\color{green}{ +1}} \\ \quad \\ \quad \\ discriminant\implies \Large {{\color{red}{ b}}^2 -4{\color{blue}{ a}}{\color{green}{ c}}}\)

OpenStudy (anonymous):

is it B

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

hmmm what discriminant did you get for B though?

OpenStudy (anonymous):

4.

OpenStudy (jdoe0001):

a = 1 b=0 c = 0 \(\bf (0)^2-4(1)(0) \implies ?\)

OpenStudy (anonymous):

oops -4

OpenStudy (jdoe0001):

hmmm -4 * 1 * 0 = ?

OpenStudy (jdoe0001):

hint: anything multiplied by 0 is 0

OpenStudy (jdoe0001):

100000 * 99999.999 * 97919849494 * one gazillion * 5 trillions * 0 = 0

OpenStudy (anonymous):

omg im sorry lol okay so 0

OpenStudy (anonymous):

whats the answer lol :/

OpenStudy (anonymous):

@jdoe0001